# Multiscale Dynamical Indices Reveal Scale-Dependent Atmospheric Dynamics

### v1

@@ Abstract
Geophysical systems are inherently complex and span multiple spatial and temporal scales, making their dynamics challenging to understand and predict. This challenge is especially pronounced for extreme events, which are primarily governed by their instantaneous properties rather than their average characteristics. Advances in dynamical systems theory, including the development of local dynamical indices such as local dimension and inverse persistence, have provided powerful tools for studying these short-lasting phenomena. However, existing applications of such indices often rely on predefined fixed spatial domains and scales, with limited discussion on the influence of spatial scales on the results. In this work, we present a novel spatially multiscale methodology that applies a sliding window method to compute dynamical indices, enabling the exploration of scale-dependent properties. Applying this framework to high-impact European summertime heatwaves, we reconcile previously different perspectives, thereby underscoring the importance of spatial scales in such analyses. Furthermore, we emphasize that our novel methodology has broad applicability to other atmospheric phenomena, as well as to other geophysical and spatio-temporal systems.
Keywords Dynamical Indices  $\cdot$ Multiscale  $\cdot$ Atmospheric Dynamics  $\cdot$ Heatwaves
@@ 1 Introduction
Since the seminar contributions by Lorenz [1, 2], it is well-know that as a result of the chaotic and multiscale nature of the atmospheric flow, the atmosphere has a limited horizon of predictability. It is also well know that such predictability is very strongly-state dependent: certain configurations of the atmospheric flow are associated with higher instability than others, so that, at practical level, the skill of weather forecast often results to be more limited [3, 4]. Lyapunov exponents (LEs) provide a powerful framework for studying the overall stability properties of a complex system and of how well we can predict its evolution on average [5]. In addition, The celebrated Kaplan-Yorke formula allows one to link the spectrum of LEs of a dynamical system to a notion of dimension of its attractor [6, 7]. Yet, again, such a notion lacks information on the local properties of the attractor. To address the local properties of the attractor, the finite-time versions of LEs - FTLEs - were introduced [8]. These convey a measure of the local (in the phase space) stability properties of the system. Indeed, in the case of the atmosphere, the local variations in the instability of the flow can be quantified by looking at the fluctuations of the individual finite-time Lyapunov exponents and at the geometry of the tangent space [9, 10, 11]. However, approaches based on LEs, including FTLEs, are more suitable for problems where an equation-based model (or its approximation) is explicitly known.
Recent advances in dynamical systems theory have provided a mathematically rigorous, and purely data-driven framework (without the need to know the underlying equation-based model of the system) for analyzing local, instantaneous, state-dependent dynamical properties of complex systems. The data-driven framework is achieved by looking at the statistics of close recurrences of the orbit with respect to a reference state of interest [12, 13]. This has led to an data-efficient way to estimate the local dimension $d$, which, roughly speaking, provides local geometric information about the system’s complexity by quantifying the number of active degrees of freedom deducible from the observations. Additionally, the analysis of recurrences has led to introducing a second index: the inverse of the persistence time ($\theta$), which estimates the reciprocal of the local persistence time near the reference state.
Given their local nature in the phase space, they are particularly well-suited for studying transient behaviors, such as extreme events in geophysical systems. Indeed, these two indices have been applied to several geophysical problems, providing insightful analysis for temperature extremes and their atmospheric drivers [14, 15, 16], the life cycles and transitions of weather regimes [17, 18, 19], ocean variability [20] and earthquake dynamics [21].
@@
The works just mentioned address the different phenomena they target looking at the system’s dynamical behavior on a single and fixed spatial domain that is of interest to the specific study. This implies that the indices adopted for each study only ‘see’ one single and fixed spatial scale, approach that does not fully capture the possibly multiscale nature of several of these and other real-world dynamics. Indeed, many geophysical and other spatio-temporal systems exhibit dynamics that span a wide range of scales, both spatially and temporally; choosing a fixed region (i.e., spatial domain) might not fully account for these multiscale physical processes. In addition, pre-selecting a fixed spatial domain requires prior knowledge of the studied system and its dynamics, and may fail to account for dynamical processes originating in regions afar from the fixed domain selected (e.g., teleconnections in the context of atmospheric dynamics).
@@
In fact, the limitations of the fixed-domain approach just outlined were also noticed by [22] in the study of tropical cyclones, where a Lagrangian perspective was proposed to better capture their dynamics. To address similar challenges, empirical mode decomposition (EMD) was recently applied to decompose a given system into components at different temporal scales, prior to calculating the local indices $d$ and $\theta$ [23, 24]. This approach reveals both the systems’ scale-dependent dynamical characteristics and the interactions between different scales [23]. However, the application of EMD to systems with multiple spatio-temporal scales remains challenging, especially if the systems are high-dimensional. At the same time, another study on the dynamics of slow earthquakes divided the spatio-temporal system into multiple subsections based on predefined physical quantities, with the dynamical properties of each subsection characterized individually [25]. Despite various efforts in the existing literature to account for the multiscale nature of real-world systems, these methods are often tailored to specific systems, thereby calling for a more systematic approach.
In this work, we seek to overcome the above-mentioned limitations by developing a spatial multiscale approach that is able to uncover dynamical properties at different spatial scales. The new approach is tailored to complex spatio-temporal systems, and based on the framework of local dynamical indices $d$ and $\theta$ [12, 13]. We show that our approach is able to provide new dynamical insights for practical real-world complex problems. To this end, we apply our approach to European summertime heatwaves, that are high-impact events that have been extensively studied in the literature, thus providing an ideal testbed for our approach [26, 27, 28, 15, 29, 30]. In particular, we show how two diverging views can be reconciled using our approach.
@@ 2 A new approach to reveal scale-dependent dynamics
Grounded at the intersection of dynamical systems theory and extreme value theory, the two instantaneous dynamical indices which we adopt here were developed by [12] and [13] to characterize local dynamical properties of complex systems. Within this framework, for each given state of interest $\zeta$, two indices—termed local dimension ($d$) and inverse persistence ($\theta$), can be computed based on the recurrences of $\zeta$ (i.e., when the system visits states neighboring $\zeta$ in the phase space). The local dimension $d$ can be considered a proxy for the number of active degrees of freedom of the system locally around $\zeta$. It therefore reflects the complexity of the dynamics near the state $\zeta$; if a system has more active degrees of freedom, its complexity is greater. The inverse persistence ($\theta$) is derived from the extremal index, which is defined in extreme value theory to measure the clustering of extremes. For a given state $\zeta$, $\theta$ can be interpreted as a measure of the inverse of the system’s persistence time near that state and is defined between 0 and 1. A high value of $\theta$ indicates that the system will quickly leave the neighborhood of $\zeta$, while a value of $\theta$ close to 0 suggests that the system will persist in states resembling $\zeta$. Both $d$ and $\theta$ can be related to the intrinsic predictability of the system at a specific time [31]. States with high local dimension ($d$) and high inverse persistence ($\theta$) are usually considered less predictable due to their complex and fast dynamics. This has been partly confirmed by [32] for both idealized and real-world datasets through a new instantaneous index that directly measures predictability within the same theoretical framework. In atmospheric applications, these dynamical indices can be computed for each instantaneous state of the atmosphere, as represented by one or more variables (e.g., a given sea-level pressure map over a given geographical domain). A more detailed description of these indices can be found in the SI Section 1, with the full mathematical derivation provided in [33]. However, it is important to acknowledge that the computation of these dynamical indices relies on assumptions that are not fully respected by the data we use, such as stationarity and hyperbolicity. The effect of this is difficult to quantify, since the framework we adopt is problematic to falsify [34].
@@
Here, we propose a novel spatially multiscale approach that computes dynamical indices using a moving window scheme, as depicted in Fig. 1. Unlike the traditional approach that considers the entire spatial domain (i.e., the entire geographical map in the case of weather applications, such as the one depicted in Fig. 1) as a state in the phase space, we treat each small map within the sliding window separately (Fig. 1a), and compute the associated dynamical indices (Fig. 1b) for each small map. Therefore, instead of obtaining one numerical value for $d$ and one for $\theta$ for the whole map, we obtain numerical values for each sliding window map for both $d$ and $\theta$. By assigning these values to the center of the sliding window map, and considering sliding windows covering the whole region of interest, we obtain spatial maps of the dynamical indices (Fig. 1c). This makes the dynamical indices not only ‘local’ in the phase space (i.e., instantaneous in time) but also spatially (i.e., geographically) ‘local’, features that can provide valuable insights. While we acknowledge that these subsystems are not strictly independent, we proceed with this division as these dynamical indices have proven effective for relatively localized systems [31].
In Fig. 1, we use a window size of 40° longitude by 20° latitude as an example, while emphasizing that the window size can be customized to capture dynamics at different scales of interest. Similarly, our use of the Euro-Atlantic region here is for illustration purposes only – our method can be extended to the entire globe, enabling the study of e.g., teleconnections in the context of weather and climate applications. It is worth noting that our framework shares structural and objective similarities with convolutional kernels from computer vision [35], which are also widely used in atmospheric dynamics studies [36, 37, 38]. Both approaches employ a sliding window technique to extract scale-dependent information from data, with one grounded in dynamical system theory and the other in machine learning.
@@ 3.1 European Summertime Heatwaves
To show the advantages of the new approach, we focus on European summertime heatwaves. These complex phenomena can pose a substantial threat to society and ecosystems [39, 40], and it is critical to enhance our understanding of their dynamics for improving prediction capabilities and mitigating their impacts [41, 42]. European summer heatwaves are driven by different processes and show significant inter-regional differences [28]. In higher-latitude regions, European summer heatwaves are often associated with atmospheric blocking, which results in adiabatic warming from subsidence and clear-sky conditions [42, 43, 30]. At lower latitudes, European summer heatwaves are linked to persistent subtropical ridges, a configuration that weakens zonal flow and strengthens meridional flow, facilitating the southerly advection of hot air into southern Europe [44]. From a traditional meteorological dynamics perspective, both blocking and subtropical ridges are highly persistent atmospheric configurations [45]. However, recent studies based on dynamical systems theory have only found a weak to moderate connection between anomalously persistent circulation patterns in the mid-troposphere and summertime heatwaves, and few significant persistence anomalies of the surface circulation patterns [15]. Meanwhile, earlier studies based on dynamical indices [13, 46], along with other research [47, 48], have argued that blocking is a transient feature of the atmospheric system, characterized by high instability and low predictability. These conclusions differ markedly from the traditional meteorological perspective and call for a reconciliation of these divergent views. In section 4, we show how our new approach can address these divergent views and enrich our dynamical understanding of these phenomena.
@@ 3.2 Data
In this study, we used state-of-the-art ERA5 reanalysis daily mean data from 1979 to 2022 [49]. The scale-dependent dynamical indices analysis is performed on both sea level pressure (SLP) and 500-hPa geopotential height (Z500), two variables extensively used to analyze circulation patterns associated with temperature extremes [15, 50, 51]. We have computed the dynamical indices for the entire Northern Hemisphere (180°W-180°E, 0° to 90°N), with the data downsampled from the original resolution of 0.25° to 0.5° to reduce computational costs. The stride size for the sliding window is set to 0.5° in both the longitude and latitude directions. In the main text, we present results using a window size of 40° longitude by 20° latitude, while the results for a smaller window size (20° longitude by 10° latitude) are provided in the Supplementary Information. Previous studies have shown that the dynamical systems metrics are dependent on the studied domain and also exhibit a strong seasonal cycle [13]. Recognizing this, we deseasonalize and standardize both dynamical indices before further interpreting the results. The seasonal cycle is calculated as the mean of a 31-day moving window centered on each calendar day, while standardization is performed on a grid-point-by-grid-point basis. For all mean values of standardized anomalies presented in this paper, we assess statistical significance using the 95% confidence interval generated by a bootstrap test with 1,000 resamples (with replacement). To control the multiple-testing issue in spatial maps, we apply the Benjamini-Hochberg procedure following [52].
We define regional summertime heatwaves over the British Isles, Scandinavia, central Europe (Germany), Iberia, and the central Mediterranean for extended summer season (JJAS), following the bounding boxes used in [28]. We then identify heatwave occurrences using the procedure outlined in [29], with the detailed steps provided in Supplementary Information.

### v2

# Spatio-temporal Dynamical Indices for Complex Systems

@@ Abstract
Complex systems span multiple spatial and temporal scales, making their dynamics challenging to understand and predict. This challenge is especially daunting when one wants to study localized and/or rare events. Advances in dynamical systems theory, including the development of state-dependent dynamical indices, namely local dimension and persistence, have provided powerful tools for studying these phenomena. However, existing applications of such indices rely on a predefined and fixed spatial domain, that provides a single scalar quantity for the entire region of interest. This aspect prevents understanding the spatially localized dynamical behavior of the system. In this work, we introduce Spatio-temporal Dynamical Indices (SDIs), that leverage the existing framework of state-dependent local dimension and persistence. SDIs are obtained via a sliding window approach, enabling the exploration of space-dependent properties in spatio-temporal data. As an example, we show that, through this framework, we are able to reconcile previously different perspectives on European summertime heatwaves. This result showcases the importance of accounting for spatial scales when performing scale-dependent dynamical analyses.
journal: Nuclear Physics B
[…]
[1]organization=National University of Singapore, country=Singapore
[…]
[2]organization=Department of Earth Sciences, Uppsala University, country=Sweden
[…]
[3]organization=Swedish centre for impacts of climate extremes (climes), Uppsala University, country=Sweden
[…]
[4]organization=Department of Meteorology and Bolin Centre for Climate Research, Stockholm University, country=Sweden
[…]
[5]organization=Laboratoire des Sciences du Climat et de l’Environnement, UMR 8212 CEA-CNRS-UVSQ, Université Paris-Saclay & IPSL, CEA Saclay l’Orme des Merisiers, city=Gif-sur-Yvette, country=France
[…]
[6]organization=Laboratoire de Météorologie Dynamique/IPSL, École Normale Supérieure, PSL Research University, Sorbonne Université, École Polytechnique, IP Paris, CNRS, city=Paris, country=France
[…]
[7]organization=London Mathematical Laboratory, city=London, country=UK
[…]
[8]organization=University of Cambridge, city=Cambridge, country=UK
[…]
[9]organization=Istituto Nazionale di Geofisica e Vulcanologia, country=Italy
[…]
[10]organization=School of Computing and Mathematical Sciences, University of Leicester, country=UK
@@ 1 Introduction
Since Lorenz’s pioneering work [1], dynamical systems theory has served as a powerful framework for studying complex systems. These systems are inherently nonlinear, high-dimensional, and multiscale, which makes them difficult to characterize and capable of exhibiting a wide range of dynamical behaviors [2, 3, 4]. While much of the existing literature has focused on the global (or average) properties of these systems [5, 6], such approaches may not be directly applicable to the study of extreme events, whose onset mechanisms are often linked to a system’s instantaneous properties. Recent advances in dynamical systems theory have provided a mathematically rigorous, and purely data-driven framework for analyzing local, instantaneous, state-dependent dynamical properties of complex systems. This data-driven framework is achieved by looking at the statistics of close recurrences of the orbit with respect to a reference state of interest $\boldsymbol{\zeta}$}, thereby inferring its dynamical properties [7, 8]. This framework provides two metrics: (i) the local dimension $d$, which is the local geometric information about the system’s complexity, and (ii) the clustering of recurrences around a state, $\theta$, which is the reciprocal of the local persistence time $\Theta=\theta^{-1}$. Indeed, these two indices have been applied to several problems, providing insightful analysis for temperature extremes and their atmospheric drivers [9, 10, 11], the life cycles and transitions of weather regimes [12, 13, 14], ocean variability [15] and earthquake dynamics [16].
@@
The works just mentioned address the different phenomena they target looking at the system’s dynamical behavior on a single and fixed spatial domain that is of interest to the specific study. This implies that the indices adopted for each study only ‘see’ one single spatial scale, an approach that does not fully capture the possibly spatially-varying dynamical nature of several of these and other real-world systems.
@@
Indeed, many geophysical and other spatio-temporal systems exhibit dynamics that span a wide range of scales, both spatially and temporally; choosing a fixed region (i.e., spatial domain) might not fully account for these multiscale physical processes. In addition, pre-selecting a fixed spatial domain requires prior knowledge of the studied system and its dynamics, and may fail to account for dynamical processes originating in regions afar from the fixed domain selected (e.g., teleconnections in the context of atmospheric dynamics) [17]. In fact, the limitations of the fixed-domain approach just outlined were also noticed by Faranda et al. [8] in the study of tropical cyclones, where a Lagrangian perspective was proposed to better capture their dynamics. To address similar challenges, empirical mode decomposition (EMD) was recently applied to decompose a given system into components at different temporal scales, prior to calculating the local indices $d$ and $\theta$ [18, 19]. This approach reveals both the systems’ scale-dependent dynamical characteristics and the interactions between different scales [18]. However, when applied to high-dimensional spatio-temporal data, the use of EMD becomes problematic, as the method does not scale efficiently and its decomposition may suffer from robustness issues in multidimensional settings [20]. Another study on the dynamics of slow earthquakes divided the spatio-temporal system into multiple subsections based on geographical locations, with the dynamical properties of each subsection characterized individually. This type of strategy can be adopted only if a priori domain-specific knowledge is available. The two examples above show that methods to characterize the multiscale nature of complex systems frequently rely on explicit decompositions or domain-specific assumptions. These examples highlight the need for a new framework that can characterize scale-dependent dynamical properties bypassing the limitations just outlined.
In this work, we address the above-mentioned limitations by introducing Spatio-temporal Dynamical Indices (SDIs), which extend the framework of local dynamical indices through the incorporation of a spatial sliding window approach. This enables a more flexible analysis of the dynamical properties of complex systems, as it no longer relies on a pre-defined domain and can be applied to uncover distinct dynamical behaviors across the spatial domain of interest. As a methodological paper, we introduce this novel approach in detail and demonstrate its effectiveness through real-world complex problem. To this end, we apply it to European summertime heatwaves – high-impact events that have been extensively studied in the literature – thus providing an ideal testbed for our method [21, 22, 23, 24, 25]. In particular, we demonstrate that our approach can reconcile two seemingly different perspectives on heatwaves — one rooted in dynamical systems theory and the other in meteorology. We note that although the scale-dependent nature of the dynamics motivates our approach, we do not systematically investigate this aspect through the use of multiple sliding windows with continuously varying sizes in this work. Instead, we base our analysis on a pre-selected window size and compare the results with those obtained over a large-scale domain. Nonetheless, we emphasize the potential of this framework to be applied across multiple spatial scales in order to obtain a more comprehensive view of the scale-dependent dynamical properties and cross-scale interactions of complex systems.
@@ 2 A new approach to reveal scale-dependent dynamics
Grounded at the intersection of dynamical systems theory and extreme value theory, the two aforementioned dynamical indices, local dimension ($d$) and inverse persistence ($\theta$) can be computed based on the recurrences of $\zeta$ (i.e., when the system visits states neighboring $\zeta$ in the phase space). The local dimension $d$ can be considered a proxy for the effective number of active degrees of freedom of the system locally around $\zeta$. It therefore reflects the complexity of the dynamics near the state $\zeta$; if a system has more active degrees of freedom, its complexity is greater. The inverse persistence ($\theta$) is derived from the extremal index, which is defined in extreme value theory to measure the clustering of extremes. For a given state $\zeta$, $\theta$ can be interpreted as a measure of the inverse of the system’s persistence time near that state and is defined between 0 and 1. A high value of $\theta$ indicates that the system will quickly leave the neighborhood of $\zeta$, while a value of $\theta$ close to 0 suggests that the system will persist in states resembling $\zeta$. Both $d$ and $\theta$ can be related to the intrinsic predictability of the system at a specific time [26]. States with high local dimension ($d$) and high inverse persistence ($\theta$) are usually considered less predictable due to their complex and fast dynamics. This has been confirmed for real-world datasets using a new instantaneous index that directly measures predictability within the same theoretical framework [27]. A more detailed description of these indices can be found in the Appendix A, with the full mathematical derivation provided in Faranda et al. [28].
When applying these dynamical indices to real-world complex systems, it is generally impossible to access all variables to construct their full phase space. Instead, we can focus on a subset of available observables (e.g., SLP for characterizing atmospheric circulation) to construct an approximated phase space. Although this is not a strict phase space in the formal sense, studying the dynamics of subsystems within this approximated phase space utilizing dynamical systems theory remains highly meaningful.
@@
In this study, we propose spatio-temporal dynamical indices (SDIs) that compute dynamical indices using a moving window scheme, as depicted in Fig. 1. In previous studies, the evolution of a variable within a fixed spatial domain (e.g., geopotential at 500 hPa, as shown in Fig. 1) has been represented as a trajectory in the approximated phase space. In contrast, we apply a sliding window with certain size across the entire domain, treating each window separately (Fig. 1a). This means that we consider the atmospheric dynamics within each window as a sub-system, for which we can construct the corresponding approximated phase space and compute its instantaneous dynamical indices (Fig. 1b). Therefore, instead of obtaining one numerical value for $d$ and one for $\theta$ for the entire domain, we obtain numerical values for each sliding window map for both $d$ and $\theta$. By assigning these values to the center of the sliding window map, and considering sliding windows covering the whole region of interest, we obtain spatial maps of the dynamical indices (Fig. 1c). This makes the dynamical indices not only ‘local’ in the phase space (i.e., instantaneous in time) but also spatially (i.e., geographically) ‘local’, features that can provide valuable insights. Our use of the Euro-Atlantic region here is for illustration purposes only – our method can be extended to the entire globe, enabling the study of e.g., teleconnections in the context of weather and climate applications.
In Fig. 1, we use a window size of 40° longitude by 20° latitude as an example, while emphasizing that the window size can be customized to capture dynamics at different scales of interest. We note that the size range of this sliding window depends on the system under study and the length of the dataset. In the case of atmospheric systems, it must be neither too small, as this would make these sub-systems overly sensitive to boundary conditions, leading to the identification of false neighbors, nor too large, which would hinder the detection of close recurrences.
Unike approaches that rely on a fixed domain [8], our spatial maps of dynamical indices capture the scale-dependent dynamical properties of complex systems, enabling a more detailed analysis of how dynamical behaviors change across spatial scales. Meanwhile, compared to previous studies that employed dynamical indices based on a Lagrangian perspective to investigate hurricanes [29], our approach can be considered a more systematic extension, with their results being a subset of ours. It is also worth noting that our framework shares structural and objective similarities with convolutional kernels from computer vision [30], a cornerstone of the field that has proven highly effective. Our method similarly employs a sliding window technique to extract scale-dependent information from data, but its foundation in dynamical systems theory enables the derivation of physically interpretable spatial maps rather than features lacking clear physical meaning.
@@ 3 A Real-World Case Study
To demonstrate the advantages of our approach, we focus on European summertime heatwaves as a case study. In this section, we briefly introduce the background and relevant research of these phenomena, as well as the data used in this study.
@@ 3.1 European Summertime Heatwaves
European summertime heatwaves can pose a substantial threat to society and ecosystems [31, 32], and it is critical to enhance our understanding of their dynamics for improving prediction capabilities and mitigating their impacts [33, 34]. These phenomena are driven by different processes and show significant inter-regional differences [23]. In higher-latitude regions, European summer heatwaves are often associated with atmospheric blocking, which results in adiabatic warming from subsidence and clear-sky conditions [34, 35, 25]. At lower latitudes, European summer heatwaves are linked to persistent subtropical ridges, a configuration that weakens zonal flow and strengthens meridional flow, facilitating the southerly advection of hot air into southern Europe [36]. From a traditional meteorological dynamics perspective, both blocking and subtropical ridges are highly persistent atmospheric configurations [37]. However, recent studies based on dynamical systems theory have only found a weak to moderate connection between anomalously persistent circulation patterns in the mid-troposphere and summertime heatwaves, and few significant persistence anomalies of the surface circulation patterns [10]. Meanwhile, earlier studies based on dynamical indices [8, 38], along with other research [39, 40], have argued that blocking is long-lived yet transient feature of the atmospheric system, characterized by anomalously high instability and low predictability. These conclusions differ markedly from the traditional meteorological perspective of blocking as a persistent atmospheric feature and call for a reconciliation of these divergent views.
In section 4, we show how our new approach can address these divergent views and enrich our dynamical understanding of these phenomena.
@@ 3.2 Data
In this study, we used state-of-the-art ERA5 reanalysis daily mean data from 1979 to 2022 [41]. The scale-dependent dynamical indices analysis is performed on both sea level pressure (SLP) and 500 hPa geopotential height (Z500), two variables extensively used to analyze circulation patterns associated with temperature extremes [10, 42, 43, 44, 45]. We remark that whilst SLP is most immediately related to describing the surface conditions of the atmosphere, the Z500 field distills the key dynamical information for the atmospheric variability of the mid-latitude because it is arguably the optimal two-dimensional field for capturing the quasi-geostrophic dynamical processes [46, 47, 48]. We have computed the dynamical indices for the entire Northern Hemisphere (180°W-180°E, 0° to 90°N), with the data downsampled from the original resolution of 0.25° to 0.5° to reduce computational costs. The stride size for the sliding window is set to 0.5° in both the longitude and latitude directions. In the main text, we present results using a window size of 40° longitude by 20° latitude, while the results for a smaller window size (20° longitude by 10° latitude) are provided in the Supplementary Information. This window size is selected as a representative scale for our analysis, as it approximately corresponds to the typical spatial extent of cyclones and anticyclones, which can be viewed as fundamental building blocks of weather regimes in the Euro-Atlantic sector [49]. We deseasonalize and standardize both dynamical indices to focus our interpretation on anomalies relative to their respective reference values.
The seasonal cycle is calculated as the mean of a 31-day moving window centered on each calendar day, while standardization is performed on a grid-point-by-grid-point basis. For all mean values of standardized anomalies presented in this paper, we assess statistical significance using the 95% confidence interval generated by a bootstrap test with 1,000 resamples (with replacement). To control the multiple-testing issue in spatial maps, we apply the Benjamini-Hochberg procedure [50].
We define regional summertime heatwaves over the British Isles, Scandinavia, central Europe (Germany), Iberia, and the central Mediterranean during the extended summer season (JJAS), following the bounding boxes used in Zschenderlein et al. [23]. We then identify heatwave occurrences using the procedure outlined in Russo et al. [24], with the detailed steps provided in Supplementary Information.